# Site-Specific Radio Channel Representation - Current State and Future Applications

### v2

@@ III-C AI/ML based channel modeling
A promising approach of SSCR is data-driven AI/ML-based channel modelling [11]. AI/ML-based channel models are trained to learn the mapping relationship between complex environmental features and channel characteristics. It can achieve a balance between generalisation and prediction accuracy compared to traditional channel models. SSCR requires the prediction of channel characteristics of a specific site according to environmental conditions, and for AI/ML-based modelling, the ability to extract features from environmental information determines the prediction accuracy of the model. For the extraction of environmental features, different data sources can be used, such as images, satellite maps, point cloud data, etc. The impact of data resolution on the accuracy of environmental features should be carefully considered. The developed AI network should take the environmental features as input and provide the corresponding site-specific channel parameters as output. Of course, physical radio propagation mechanisms and empirical channel models can be used and incorporated into the designed AI model to improve the site-specific channel prediction. Methods such as a regularisation strategy, an early stop strategy and ensemble learning can also be used to avoid model over-fitting for a specific scenario. In addition, the ability of the model to self-evolve for different sites should be used as an evaluation index, and its essence lies in strengthening the existing model with additional data and features. Sufficient multi-source data and prior knowledge can support AI/ML-based channel modelling for SSCR needs with low complexity.
IV Model validation
IV-A Non-stationary environment evaluation
@@
The local scattering function (LSF) [3] can be used to evaluate non-stationary measurement data. With the time-variant frequency response estimate , the LSF can be computed as
@@
with the Doppler index , the delay index , and the tapers are two-dimensional discrete prolate spheroidal sequences as shown in detail in [3]. The number of tapers controls the bias-variance trade-off of the LSF estimate. Using the LSF we can compute the time-variant power delay profile (PDP), Doppler spectral density (DSD) and path loss as marginals with respect to Doppler, time or both.
IV-B Frequency domain continuity
@@
Another validation requirement for SSCR is frequency domain continuity, i.e. ensuring the applicability of channel parameters over the entire frequency range (FR) of operation. 5G NR pushes the traditional sub- GHz boundary to GHz (FR) and introduces millimeter wave (mmWave) bands from GHz to GHz (FR). The range in between, from GHz to GHz, is called the upper mid-band (FR) and is being considered for 5G-Advanced and 6G. Experiments in the sub-THz range from - GHz (FR) and - GHz (FR), as well as in the THz range ( GHz) have attracted interest. The validation of parameters over such different frequency ranges prompted a series of simultaneous multi-band measurements. The blocking and deviation characteristics of a scatterer become more pronounced as the carrier frequency increases. This means that small scatterers, which could be safely ignored for FR or FR bands, may need to be included in the SSCR when applied to bands for FR or above.
IV-C Joint communication and sensing
@@
Radar cross-section parameters need to be included for path loss calculation in monostatic/bistatic sensing. If the link lengths for TX-target and target-RX are given by and , then the path loss is given by , where is the normalised reflectivity of the target. Of course, measuring the radar cross section of different targets and classifying them according to their radar cross section range are important objectives for JCAS field test campaigns. Additional SSCR validation parameters include clutter/scattering patterns and object mobility, depending on the accuracy, resolution, and refresh rate requirements of the JCAS use case. The return power from natural or man made backgrounds is also critical for defining the target to noise power ratio.
IV-D Empirical radio channel measurements
@@
In earlier models pathloss, delay spread and Doppler spread were modeled as function of the environment. SSCRs aim to include also their dependence on distance as well as on the surrounding environment and its geometrical properties. As an example we depict the root mean square (RMS) delay spread cumulative distribution function (CDF) for two medium scale urban routes that have different orientations with respect to the TX (in Manchester, UK, at 2.1 GHz) in Fig. 5.
New 5G and beyond inspired use cases, as described in Section II, require new empirical radio channel measurements for verification:
IV-D1 Vehicular scenarios
For rail and road applications, new measurement campaigns are aimed at capturing the interplay between higher centre frequency and highly mobile site-specific features. Challenges for such campaigns include ensuring repeatable and fault-free operation of the channel sounder under extreme environmental conditions (weather, vibration, etc.) and maintaining proper synchronization between different subsystems. The node mobility causes fluctuations in the beamwidth, and as a result, the number of MPCs varies proportionally with the beamwidth. The channel becomes highly non-stationary, and a local scattering function, as described in subsection IV-A, can be used to characterize the channel.
IV-D2 Airborne scenarios
For airborne scenarios, the mobile-to-mobile (M2M) channel model validation requires at least two additional facets: first, consideration of the full three-dimensional site geometry (unlike vehicular scenarios where a 2D model is generally sufficient), and second, consideration of scattering components (along with line-of-sight and specular reflections). The validation is done in the delay-Doppler plane to account for the TX and RX’s fast and unpredictable mobility patterns. Classical site-general parameters, such as shadowing variance, are augmented with new features like shadowing correlation distances to ensure the validity of an airborne M2M channel model at different TX/RX heights.
IV-D3 Urban cellular high frequency scenarios
For example, [14] builds on the quasi deterministic channel model type described in Sec. III-B. The authors present a recipe for mmWave MPC generation for a specific urban cellular environment in Yokohama, Japan. Here, the measured data are used to estimate an exponential decay model to calibrate the ray tracer. Random clusters are generated with the measured site-specific statistical parameters of large and small scale fading parameters. This step improves the accuracy of the quasi-deterministic channel model by incorporating real-world measurement data that captures the specific characteristics of the environment under consideration.
V Channel emulation
@@
In [12] a novel site-specific channel emulation approach is presented where the convolution is approximated with a reduced rank subspace representation of dimension using discrete prolate spheroidal sequences. This approach avoids the quadratic increase of with and enables site-specific channel emulation. In Fig. 4 we reproduce [12, Fig. 13c], showing the time-variant PDP from the sub-space based radio channel emulator that is controlled by an SSCR. A good match with the measured radio propagation scenario in [12, Fig. 13b] is achieved.
@@
VI Conclusions and outlook
In this paper, we have reviewed the motivations for the inclusion of SSCR in future standards (e.g. 5G and beyond and WiFi7) and introduced the main use cases such as broadband communication with users in large buildings, JCAS, and reliable communication links with railways, road vehicles and aircraft. Ray tracing, quasi-deterministic and AI/ML-based models have been introduced and their specific advantages and disadvantages for SSCRs have been explored. Model validation methods for non-stationary environments, JCAS and frequency domain continuity requirements are explained. Empirical radio channel measurements for validation of path loss characteristics, vehicular, airborne and urban cellular scenarios are reviewed. Finally, low-complexity channel emulation techniques for site-specific hardware-in-the-loop testing of communications hardware are presented.
SSCRs is an emerging field with great importance for future wireless communication systems that will utilise D-MIMO, RIS, multi-band communication, and JCAS techniques. The combination of an SSCR with advanced PHY layer technologies will enable reliable and energy efficient communication systems of the future.

### v4

@@ V-C AI/ML based method
The developed AI network should take the environmental features as input and provide the corresponding site-specific channel parameters as output. Of course, physical radio propagation mechanisms and empirical channel models can be used and incorporated into the designed AI model to improve the site-specific channel prediction. Methods such as a regularization strategy, an early stop strategy and ensemble learning can also be used to avoid model over-fitting for a specific scenario. In addition, the ability of the model to self-evolve for different sites should be used as an evaluation index, and its essence lies in strengthening the existing model with additional data and features. Sufficient multi-source data and prior knowledge can support AI/ML-based channel modeling for SSCR needs with low complexity.
The combined use of the three methods presented above is a promising future research direction to achieve the best trade-off between complexity and accuracy.
VI Step by step description for SSCR setup and validation
Here we describe the steps needed to setup a SSCR and use it for numerical link level simulation or for real-time channel emulation. The description illustrates the procedure for the quasi deterministic method for a vehicular scenario in detail but also indicates the basic procedure for the two other methods and other scenarios.
1.
Environment definition: Obtain a basic environment geometry from a database such as, e.g. open street map. Augment map data with objects detected via LIDAR and video data. Figure 3 shows an exemplary quasi-deterministic environment model for a vehicular scenario including buildings, as well as additional vegetation, parked cars, and city lamp posts. This additional objects improve the channel modeling accuracy substantially in terms of time-variant PDP and time-variant DSD as depicted in [11, Fig. 5].
Figure 3: Quasi-deterministic environment model for a vehicular scenario, see [11, Fig. 4b]. Red dots indicate point scatterers randomly distributed along walls and vegetation to represent static diffuse scattering, mobile discrete scatterers are represented by blue squares and circles, black squares represent static discrete scatterers such as parked cars or street lamps. -
2.
Evaluate:
Ray tracing: Use geometrical optics to compute line-of-sight, reflection up to a specified order , diffraction and diffuses scattering MPCs.
Quasi deterministic: Distribute point scatterers in the geometry domain (shown as red dots in Fig. 3) along surfaces of objects according to a given distribution and compute MPCs between TX and RX taking blocking by buildings into account (alternatively scatterers can be also placed in the delay/angle domain). Use the MPC path weight to limit the number of relevant MPCs for each stationarity region.
AI/ML: (i) Train the AI/ML model with environment data and measured channel transfer functions for given TX and RX coordinates. (ii) Infer MPCs from the AI/ML model based on new coordinates of TX and RX.
- 3.
4.
Real-time emulation: Use the channel model for real-time emulation employing a hardware-in-the-loop (HiL) setup, see Section IX for more details. Real-time validation can be performed, e.g., comparing the frame error rate obtained in the HiL setup with the one obtained by empirical measurements. In Fig. 4 (reproduced from [11, Fig. 6]) the frame error rate (FER) measured on the road is compared with the FER obtained by emulating the channel transfer function from the quasi deterministic SSCR. Two cases are shown: (i) the case when all identified objects are included, and (ii) the case when only objects from open street map are included. A close match in terms of FER between measurement and SSCR based emulation can be observed in Fig. 4 for the former case (i).
Figure 4: FER comparison reproduced from [11, Fig. 6]: FER obtained directly on the road, FER obtained by emulating channel transfer function from the GSCM in the case when all identified objects are included, and FER obtained by emulating CTF from the GSCM in the case when only objects from open street map are included. The violet region indicates the 95% confidence interval (CI) of the binomially distributed FER.
VII Model validation
An SSCR shall be able to represent non-stationary fading conditions while being frequency and spatially (i.e. time) consistent for the evaluation of 5G and 6G technologies that use large antenna apertures, perform multi-band operation and implement JCAS algorithms. To validate an SSCR we need to establish respective validation methods by opening new research directions, such as:
VII-1 Non-stationary environment evaluation
@@
The local scattering function (LSF) [3] can be used to evaluate non-stationary measurement data and compare it with the SSCR outcome. With the time-variant frequency response estimate , the LSF can be computed as
@@
with the Doppler index , the delay index , and the tapers are two-dimensional discrete prolate spheroidal sequences as shown in detail in [3]. The number of tapers controls the bias-variance trade-off of the LSF estimate. Using the LSF we can compute the time-variant PDP, DSD and path loss as marginals with respect to Doppler, time or both.
VII-2 Frequency domain continuity
@@
Another validation requirement for SSCR is frequency domain continuity, i.e. ensuring the applicability of channel parameters over the entire frequency range of operation. 5G NR pushes the traditional sub- GHz boundary to GHz (FR) and introduces millimeter wave (mmWave) bands in FR. FR is called the upper mid-band and is being considered for 5G-Advanced and 6G. Experiments in the sub-THz range from - GHz (FR) and - GHz (FR), as well as in the THz range ( GHz) have attracted interest. The validation of parameters over such different frequency ranges prompted a series of simultaneous multi-band measurements. The blocking and deviation characteristics of a scatterer become more pronounced as the carrier frequency increases. This means that small scatterers, which could be safely ignored for FR or FR bands, may need to be included in the SSCR when applied to bands for FR or above.
VII-3 Backscatter radio channel properties
@@
JCAS requires new radio channel properties to be represented by an SSCR. Hence, radar cross-section parameters need to be included for path loss calculation in mono-/bi-static sensing. If the link lengths for TX-target and target-RX are given by and , then the path loss is given by , where is the normalized reflectivity of the target. Measuring the radar cross section of different targets and classifying them according to their radar cross section range are important objectives for JCAS field test campaigns. Additional SSCR validation parameters include clutter/scattering patterns and object mobility, depending on the accuracy, resolution, and refresh rate requirements of the JCAS use case. The return power from natural or man made backgrounds is also critical for defining the target to noise power ratio.
VIII Empirical radio channel measurements
@@
In earlier models pathloss, delay spread and Doppler spread were modeled as function of the environment. SSCRs aim to include also their dependence on distance as well as on the surrounding environment and its geometrical properties. As an example we depict the root mean square (RMS) delay spread cumulative distribution function for two medium scale urban routes that have different orientations with respect to the TX (in Manchester, UK, at 2.1 GHz) in Fig. 5.
New 5G and 6G inspired use cases, as described in Section II, require new empirical radio channel measurements for verification, which constitute interesting future research directions:
VIII-1 Vehicular scenarios
For rail and road applications, new measurement campaigns are aimed at capturing the interplay between higher centre frequency and highly mobile site-specific features. Challenges for such campaigns include ensuring repeatable and fault-free operation of the channel sounder under extreme environmental conditions (weather, vibration, etc.) and maintaining proper synchronization between different subsystems. The node mobility causes fluctuations in the beamwidth, and as a result, the number of MPCs varies proportionally with the beamwidth. The channel becomes highly non-stationary, and a local scattering function, as described in subsection VII-1 and [3], can be used to characterize and compare the measured and modelled channel.
VIII-2 Airborne scenarios
For airborne scenarios, the mobile-to-mobile (M2M) channel model validation requires at least two additional facets: first, consideration of the full three-dimensional site geometry (unlike vehicular scenarios where a 2D model is generally sufficient), and second, consideration of scattering components (along with line-of-sight and specular reflections). The validation is done in the delay-Doppler plane to account for the fast mobility patterns of TX and RX. Classical site-general parameters, such as shadowing variance, are augmented with new features like shadowing correlation distances to ensure the validity of an airborne mobile-to-mobile channel model at different TX/RX heights.
VIII-3 Urban cellular high frequency scenarios
For example, [13] builds on the quasi deterministic channel model type described in Sec. V-B. The authors present a recipe for mmWave MPC generation for a specific urban cellular environment in Yokohama, Japan. Here, the measured data is used to estimate an exponential decay model to calibrate the ray tracer. Random clusters are generated with the measured site-specific statistical parameters of large and small scale fading parameters. This step improves the accuracy of the quasi-deterministic channel model by incorporating real-world measurement data that captures the specific characteristics of the environment under consideration.
IX Channel emulation
@@
In [14] a novel site-specific channel emulation approach is presented where the convolution is approximated with a reduced rank subspace representation of dimension using discrete prolate spheroidal sequences. This approach that is fundamentally MPC based, avoids the quadratic increase of with and enables site-specific channel emulation. In Fig. 6 we reproduce [14, Fig. 13c], showing the time-variant PDP from the subspace based radio channel emulator that is controlled by an SSCR. A good match with the measured radio propagation scenario in [14, Fig. 13b] is achieved.
Combining the subspace based channel emulation methods with a quasi deterministic SSCR allows for real-time hardware-in-the-loop emulation [11].
@@
X Conclusions and future research
In this paper we have reviewed the motivations for including SSCRs in future standards (e.g. 5G, 6G and WiFi 7). Ray tracing, quasi-deterministic and AI/ML-based methods were introduced and their specific advantages and disadvantages for SSCRs were explored. Model validation methods for non-stationary environments, frequency domain continuity, and backscatter radio channel characteristics were explained. Empirical radio channel measurements for validation of path loss characteristics, urban cellular, vehicular, and airborne scenarios are reviewed. Finally, low-complexity channel emulation techniques for hardware-in-the-loop testing of communications hardware using an SSCR are presented.
Future research directions for SSCRs are (a) acquisition of empirical measurement data for vehicular and airborne scenarios, (b) methods to compute MPCs that achieve a good tradeoff between accuracy and complexity, (c) new model validation techniques, and (d) real-time channel emulation methods that directly exploit the MPC-based architecture of SSCRs.
SSCRs are an emerging field with great importance for 5G and 6G wireless communication systems to support research and development of emerging D-MIMO, RIS, multi-band communication, and JCAS techniques.
Acknowledgement
We would like to thank COST INTERACT for helpful discussion.